solve f(x, u_x,u_z)=g(y,u_y,u_z) by Jacobi method.
1
Expert's answer
2020-07-05T17:34:22-0400

Let "f(x,y,z,u_x,u_y,u_z) = 0" (1)

The Fundamental idea of the Jacobi method is to introduce two first order PDE involving two arbitrary constants a and b of the following form,

"h_1(x,y,z,u_x,u_y,u_z,a) = 0" (2)

"h_2(x,y,z,u_x,u_y,u_z,b) = 0" (3)

such that "\frac{\partial(f,h_1,h_2)}{\partial(u_x,u_y,u_z)} \neq 0" (4)


and

Equations (1),(2) and (3) can be solved for "u_x,u_y,u_z"

and equation "du = u_xdx + u_ydy+u_zdz" is integrable


If "h_1 = 0" and "h_2=0" are compitable with"f =0" then "h_1,h_2" satisfy,

"\frac{\partial(f,h)}{\partial(x,u_x)}+\frac{\partial(f,h)}{\partial(y,u_y)}+\frac{\partial(f,h)}{\partial(z,u_z)} =0" for "h = h_i", i =1,2

Last equation will lead to the PDE of the form,

"f_{u_x}\frac{\partial h}{\partial x}+f_{u_y}\frac{\partial h}{\partial y}+f_{u_z}\frac{\partial h}{\partial z} - f_x\frac{\partial h}{\partial u_x}- f_y\frac{\partial h}{\partial u_y}- f_z\frac{\partial h}{\partial u_z} = 0"

It's auxiliary equation will be,

"\frac{dx}{f_{u_x}}=\frac{dy}{f_{u_y}}=\frac{dz}{f_{u_z}} = \frac{du_x}{-f_x}= \frac{du_y}{-f_y}= \frac{du_z}{-f_z}"

After getting this we need to those fractions which will lead to the solution.


Since given equation is

"f(x,u_x,u_z) = g(y,u_y,u_z)"

it can be simplified as

"f(x,u_x,u_z) - g(y,u_y,u_z) = 0"

or it can be written as

"F(x,y,u_x,u_y,u_z) = 0"

Then auxiliary equation will be

"\frac{dx}{F_{u_x}}=\frac{dy}{F_{u_y}} =\frac{dx}{F_{u_z}}= \frac{du_x}{-F_x}= \frac{du_y}{-F_y}"

Solve it to obtain the solution.